# Conceptualizing Uncertainty

### v1

@@ 4.1 Rejecting Uncertain Points with Concepts
Figure 3: (left) Accuracy-Rejection ($\uparrow$) and (right) OOD-Rejection Curves ($\downarrow$).
We demonstrate that our proposed explanations enable actionable interpretations by utilizing the learned concepts for rejection decisions in a classification setting. In particular, given a trained model $\mathbf{M}$ and a set of new data points $\mathbf{X}$, we apply our proposed approach to generate concept banks $\mathbf{V}_{\text{CER}},\mathbf{V}_{\text{UNC}}$, according local importances $e_{l}(\mathbf{x}_{i}),\forall\mathbf{x}_{i}\in\mathbf{X}$ and global ones $e_{\text{CER}},e_{\text{UNC}}$.
Now we consider the question: Can we improve a basic rejection strategy, which rejects a fixed amount of points based on predictive uncertainty, by leveraging our explanations? For the latter, we consider two strategies: 1. Concept-only rejection: we identify for each input $\mathbf{x}_{i}$ the most strongly activated concept $c^{*}=\arg\max\mathbf{W}_{i}$ utilizing the combined concept bank $[\mathbf{V}_{\text{CER}},\mathbf{V}_{\text{UNC}}]$ and determine the global importance of $c^{*}$. We then reject those points first, for which $c^{*}\in e_{\text{UNC}}$ and with highest $e_{\text{UNC}}$ value. This leads to inputs associated with globally uncertain concepts being rejected first, while those linked to globally certain concepts being retained longer. 2. Weighted rejection: We adapt the previous strategy, by weighting the uncertainty output $f(\mathbf{x}_{i})$ with $+1$ or $-1$, depending on whether $c^{*}\in e_{\text{UNC}}$ or $c^{*}\in e_{\text{CER}}$ and again reject according to highest value.
To evaluate the effectiveness of these methods, we use a pre-trained ResNet-50 classifier as the base model. We randomly sample images from 10 out of 20 ImageNet classes (ImageWoof + Imagenette) and also include out-of-distribution (OOD) samples from the NINCO dataset [3]. In each run, we use 1,000 points, with 40% being OOD to compute accuracy-rejection curves [23, 25], which depict the accuracy computed on all not-rejected points on the y-axis for different percentage rejection rates (x-axis). We compare our two concept-based strategies, Concept-only and Weighted rejects, to baselines, which reject according to highest predictive uncertainty, using each of $\hat{u}_{t},\hat{u}_{e},\hat{u}_{a}$. We also compute the AUC for each strategy and average these measures over 20 runs. We set the number of concepts to 55 for the certain group and 35 for the uncertain.
As seen in Figure 3 (left), our "weighted" method performs with the highest AUC score computed across 20 runs. To confirm the statistical significance of this result, we conduct a one-sided Wilcoxon signed rank test against the Total Entropy (which in this case is very similar to the Aleatoric curve), obtaining a p-value $<10^{-6}$, indicating ours performing significantly better. Meanwhile, the concept-only strategy performs better for medium rejection rates and worse for higher, as compared to the baselines.
In Figure 3 (right), we plot the percentage of OOD points among the not-rejected ones. Up to approximately 20% rejection, all curves except the Epistemic and Concept-only one, exhibit similar performance. However, beyond this point, both our strategies reject a greater proportion of OOD samples at each step. This signifies why our explanations improve the rejection: the learned concepts pick up on the OOD data and helps in rejecting them.
4.2 Distinguishing Sources for Uncertainty
Figure 4: (left) Motion Blur experiment results ($\uparrow$) and an example noisy image (right) to better visualize the noise and the corresponding concept extracted.
In particular, our setting involves a classifier $\mathbf{M}$ that is applied in the wild and is exposed to two different types of noise. We consider structured noise induced by novel classes not seen during training and unstructured noise, like random distortions (e.g., blurring). We want to evaluate in how far our explanations can help to detect the different noise types and filter the instances with unstructured noise out, without having explicit knowledge about the noise beforehand. This can occur in an active learning setup, where a data exploration task could be to filter out unrecognizable noisy images that provide no meaningful learning signal and finetune the model on the remainder.
More precisely, we consider the setting where new data samples $\mathbf{X}$ are available and our explanation pipeline is applied to provide concept banks $\mathbf{V}_{\text{CER}},\mathbf{V}_{\text{UNC}}$, according local importances $e_{l}(\mathbf{x}_{i}),\forall\mathbf{x}_{i}\in\mathbf{X}$ and global ones $e_{\text{CER}},e_{\text{UNC}}$. We aim to evaluate how well $\mathbf{V}_{\text{UNC}}$ captures different noise types, and how well unstructured noise samples $\mathbf{x}_{i}$ can be filtered out using $e_{l}(\mathbf{x}_{i})$ and $\mathbf{W}_{i}$ of $\mathbf{V}_{\text{UNC}}$ concepts. For this purpose, we assume that we can determine the set $nc$ that correspond to unstructured noise concepts in $\mathbf{V}_{\text{UNC}}$ (as a human could visually inspect them). Then, for each data point $\mathbf{x}_{i}$, we sum the local importance $e_{l,nc}(\mathbf{x}_{i})$ or NMF activations $\mathbf{W}_{i,nc}$ and filter out according to highest values, corresponding to the amount of presence of these concepts. We refer to these strategies as Ours (Importance) when using $e_{l,nc}(\mathbf{x}_{i})$, and to and Ours (NMF) for $\mathbf{W}_{i,nc}$.
@@ 4.2 Distinguishing Sources for Uncertainty
Table 1: We report the average AUC score ()($\uparrow$)( ↑ ) over 20 runs for various types of random noise patterns. Our methods using the local importances or reduced NMF representation perform better than the baselines. We test our methods against Gaussian Blurring (G Blur), Salt and Pepper Noise (S and P), Gaussian Noise (G Noise), Motion Blurring (M Blur), Radial Blurring (R Blur), and Wave Noise(Wave)
Method Total Aleatoric Epistemic Ours (Imp) Ours (NMF)
G Blur $79.6\pm 1.5$ $79.5\pm 1.5$ $82.6\pm 1.2$ $\underline{85.6}\pm 1.3$ $\textbf{89.0}\pm 0.2$
SnP Noise $77.2\pm 1.4$ $77.3\pm 1.4$ $76.7\pm 1.7$ $\underline{85.4}\pm 0.9$ $\textbf{88.9}\pm 0.2$
G Noise $82.3\pm 1.3$ $82.5\pm 1.3$ $70.8\pm 2.7$ $\underline{85.1}\pm 1.2$ $\textbf{88.8}\pm 0.3$
M Blur $77.5\pm 2.0$ $77.4\pm 2.0$ $80.4\pm 1.4$ $\underline{87.0}\pm 0.9$ $\textbf{89.2}\pm 0.2$
R Blur $82.2\pm 1.6$ $82.4\pm 1.6$ $73.1\pm 2.9$ $\underline{85.2}\pm 4.4$ $\textbf{86.2}\pm 4.9$
Wave Noise $80.8\pm 7.4$ $80.1\pm 8.0$ $87.8\pm 2.0$ $\textbf{88.3}\pm 1.5$ $\textbf{88.3}\pm 1.5$
Average $79.9\pm 2.2$ $79.9\pm 2.3$ $78.5\pm 6.3$ $\underline{86.1}\pm 1.3$ $\textbf{88.4}\pm 1.1$
The results in Table 1 indicate that our method outperforms uncertainty-based measures for this task. Thereby, using $e_{l}(g(\mathbf{x}_{i}))$ achieves the second-best performance (underlined), while leveraging $\mathbf{W}_{i}$ yields the best results (bolded), with the highest average AUC score. Inspecting Figure 4, we can see that our method using the NMF coefficients (purple) consistently recommends images that would be beneficial to the re-training set while abstaining from recommending the unusable images until they must be chosen. We indicate the true unusable image percentage by the vertical line on the figure. In the ideal case, a method would not recommend any unusable images until it reaches the vertical line. Our method provides an explainable way to automatically select images which contain meaningful structure for domain adaptation.
@@ 4.3 Explaining Uncertainty in Language Models for Fairness
For this purpose, we fine-tune a BERT [10] model on the Bias in Bios dataset [7] which consists of biographies where the task is to predict their corresponding occupation. Before fine-tuning, we incorporate ReLU on the last embedding layer to ensure non-negativity. We then apply our pipeline using a NMF-based concept extraction technique for the text domain [18]. We inspect the "physician" class, by training an NMF on the inputs predicted as such and compute their importances with respect to each group of uncertain and certain points.
We investigate the most important concept of the uncertain samples, number 6, in the following. In Figure 5, to understand what concept 6 represents, we show two excerpts which activate concept 6. Thereby, the intensity of red marks the strength of the activation of concept 6. We can see that female pronouns appear among the highlighted words along with other nouns like "co-founder" and "sleep medicine specialist". Since we know the gender labels of the dataset, we check the Pearson correlation between concept 6 and the labels. Indeed, it is the most correlated with gender at $R=0.3$. We further verify the relevance of this concept for representing gender by excluding it in the NMF reconstruction and applying the occupation classifier. This changes some of the predictions, most notably a large proportion of professors and chiropractors that were falsely predicted as physicians. Even more interesting, the gender distribution among the samples influenced by concept 6 does not align with the gender distribution of said classes, which could be an indication of gender bias in BERT. We evaluate the change in gender bias by computing the equalized odd score before and after our intervention and report an improvement of 0.0027. At first glance this score does not sound impressive, the intervention on the physician class only affects a limited number of samples and thus has a limited effect on global equalized odds. More precisely, the best possible outcome for an intervention on the physician class (fixing all false positives) would have led to an improvement of 0.0069. This demonstrates that the detected relevant concept in the uncertain group encodes gender information in our example, and that removing it can improve fairness.
@@ 5 Limitations and Future Work
While our approach performs well in the tasks outlined above, it is not without limitations. Concept-based explanations provide a human-interpretable means of understanding uncertainty in machine learning settings, particularly when the source of uncertainty is visibly discernible. However, they may fail to capture finer-grained pixel-level nuances of uncertainty. In this study, we maintained a fixed patch size when training the NMF, but varying the patch size could potentially capture more localized properties of uncertainty. Investigating the relationship between patch size and its impact on concept-based explanation quality presents a promising direction for future research.
Additionally, our method for estimating concept importance may not be optimal, as evidenced by the superior performance of using NMF activations directly in our experiments. We suspect this performance gap arises from the variance—or lack thereof—of the uncertainty measure. Specifically, if we perturb a concept within an already highly uncertain input, the uncertainty measure may not exhibit significant variation. While we did not explore alternative ways to refine importance attribution in this study, we do plan to do address it in future iterations of this work.
@@ 6 Conclusion
We introduced a novel framework for explaining uncertainty using automatically extracted concept activation vectors. Our proposed framework enables both local and global explanations of uncertainty through the use of importance scores and attribution maps. These explanations demonstrate their utility by being incorporated into rejection strategies, aiding design of useful re-training sets, and helping to detect and mitigate bias. Moreover, while concept-based explanations of model predictions can be useful, using CAVs to capture sources of uncertainty not only offers another complementary view into how a model makes its decisions, but also provides interpretable ways to enhance its performance.

### v2

@@ 4.2 Rejecting Uncertain Points with Concepts
We demonstrate that our proposed explanations encapsulate uncertainty by utilizing the learned concepts in a classification setting to improve decision-making. Often, uncertainty estimations are evaluated indirectly by measuring the improvement of predictions [19] through accuracy-rejection curves, which depict the accuracy (y-axis) of a classifier as a function of its rejection rates (x-axis) [25, 33]. If the estimation performs well, we should expect the curve to be monotonically increasing. In this experiment, we create a rejection strategy using our explanations and thus evaluate their effectiveness as an uncertainty estimator. We also compare the results to baseline uncertainty estimations. In particular, given a trained model $\mathbf{M}$ and a set of new data points $\mathbf{X}$, we apply our proposed approach to generate concept banks $\mathbf{V}_{\text{CER}},\mathbf{V}_{\text{UNC}}$, according local importances $e_{l}(\mathbf{x}_{i}),\forall\mathbf{x}_{i}\in\mathbf{X}$ and global ones $e_{\text{CER}},e_{\text{UNC}}$.
Now, we build two strategies: 1. Concept-only rejection: we identify for each input $\mathbf{x}_{i}$ the most strongly activated concept $c^{*}=\arg\max\mathbf{W}_{i}$ utilizing the combined concept bank $[\mathbf{V}_{\text{CER}},\mathbf{V}_{\text{UNC}}]$ and determine the global importance of $c^{*}$. We then reject those points first, for which $c^{*}\in e_{\text{UNC}}$ and with the highest $e_{\text{UNC}}$ value. This leads to inputs associated with globally uncertain concepts being rejected first, while those linked to globally certain concepts are retained longer. 2. Weighted rejection: We adapt the previous strategy, by weighting the uncertainty output $f(\mathbf{x}_{i})$ with +1+1+ 1 or 1-1- 1, depending on whether $c^{*}\in e_{\text{UNC}}$ or $c^{*}\in e_{\text{CER}}$ and again reject according to highest value.
Concerning the implementation, we use a pre-trained ResNet-50 classifier as the base model. We randomly sample images from 10 out of 20 ImageNet classes (ImageWoof [18] + Imagenette [18]) and also include out-of-distribution (OOD) samples from the NINCO dataset [3]. In each run, we use 1,000 points, with 40% being OOD to compute the accuracy-rejection curves. We additionally compare our two concept-based strategies, Concept-only and Weighted rejects, to baselines, which reject according to highest predictive uncertainty, using each of $\hat{u}_{t},\hat{u}_{e},\hat{u}_{a}$. We also compute the AUC for each strategy and average these measures over 20 runs. We set the number of concepts to 55 for the certain group and 35 for the uncertain.
As seen in Figure 5 (left), our “weighted” method is a monotonically increasing curve and performs with the highest AUC score computed across 20 runs. To confirm the statistical significance of the AUC score, we conduct a one-sided Wilcoxon signed rank test against the Total Entropy (which in this case is very similar to the Aleatoric curve), obtaining a p-value $<10^{-6}$, indicating ours performing significantly better. Meanwhile, the “concept-only” strategy is also monotonically increasing and performs better for medium rejection rates and worse for higher ones, in comparison to the baselines.
In Figure 5 (right), we plot the percentage of OOD points as a function of the rejection rate. Up to approximately 20% rejection, all curves except the Epistemic and Concept-only one, exhibit similar performance. However, beyond this point, both our strategies reject a greater proportion of OOD samples at each step. This signifies why our explanations work well as an uncertainty estimator and better than the baselines: the learned concepts pick up on the OOD data and aid in their rejection.
@@ 4.3 Explaining Uncertainty in Language Models for Fairness
For this purpose, we fine-tune a BERT [10] model on the Bias in Bios dataset [7] which consists of biographies where the task is to predict their corresponding occupation. Before fine-tuning, we incorporate ReLU on the last embedding layer to ensure non-negativity. We then apply our pipeline using a NMF-based concept extraction technique for the text domain [20]. We inspect the "physician" class by training an NMF on the inputs predicted as such and compute their importances with respect to each group of uncertain and certain points. We investigate the most important concept of the uncertain samples, concept 6, in the following analysis.
In Figure 6, to understand what concept 6 represents, we show two excerpts that activate concept 6. Thereby, the intensity of red marks the strength of the activation of concept 6. We can see that female pronouns appear among the highlighted words along with other nouns like "co-founder" and "sleep medicine specialist". Since we know the gender labels of the dataset, we check the Pearson correlation between concept 6 and the labels. Indeed, it is the most correlated with gender at $R=0.3$. We further verify the relevance of this concept for representing gender by excluding it in the NMF reconstruction and applying the occupation classifier. This changes some of the predictions, most notably a large proportion of professors and chiropractors who were falsely predicted as physicians. Even more interesting, the gender distribution among the samples influenced by concept 6 does not align with the gender distribution of said classes, which could be an indication of gender bias in BERT. We evaluate the change in gender bias by computing the equalized odds score before and after our intervention and report an improvement of 0.0027. At first glance, this score does not sound impressive, but it is worth noting that the intervention on the physician class only affects a limited number of samples and thus has a limited effect on global equalized odds. More precisely, the best possible outcome for an intervention on the physician class (fixing all false positives) would have led to an improvement of 0.0069. This demonstrates that the detected relevant concept in the uncertain group encodes gender information in our example and that removing it can improve fairness.
@@ 5 Limitations and Future Work
While our approach performs well in the tasks outlined above, it is not without limitations. Concept-based explanations provide a human-interpretable means of understanding uncertainty in machine learning settings, particularly when the source of uncertainty is visibly discernible. However, they may fail to capture finer-grained pixel-level nuances of uncertainty. In this study, we maintained a fixed patch size when training the NMF, but varying the patch size could potentially capture more localized properties of uncertainty. Investigating the relationship between patch size and its impact on concept-based explanation quality presents a promising direction for future research. This limitation is further exemplified by changing the OOD percentage in Experiment 4.2 from 40% to 20% as seen in the Appendix. We observe the convergence of the performance of our methods with the baselines. The OOD data provides clear differences captured by concepts, such that we can reject a classification based on an input’s most activated concept; however, when the number of OOD is decreased, the source of the uncertainty becomes more subtle between known classes that are difficult to discriminate.
Additionally, our method for estimating concept importance may not be optimal, as evidenced by the superior performance of using NMF activations directly in our experiments. We suspect this performance gap arises from the variance, or lack thereof, of the uncertainty measure. Specifically, if we perturb a concept within an already highly uncertain input, the uncertainty measure may not exhibit significant variation. While we did not explore alternative ways to refine importance attribution in this study, we do plan to address it in future work.
Finally, while the usefulness of our concept-based uncertainty explanations is evaluated in downstream tasks, they do elicit downstream human action and decision-making. This prompts a user-centric study to evaluate the effectiveness of our proposed explanations in explaining uncertainty to a user. Such a study is a subject for future work.
@@ 6 Conclusion
We introduced a novel framework for explaining uncertainty using automatically extracted concept activation vectors. Our proposed framework enables both local and global explanations of uncertainty through the use of importance scores and attribution maps. These explanations demonstrate their utility by encapsulating uncertainty, aiding the design of useful re-training sets, incorporating them into rejection strategies, and helping to detect and mitigate bias. Moreover, while concept-based explanations of model predictions can be useful, using CAVs to capture sources of uncertainty not only offers another complementary view into how a model makes its decisions but also provides interpretable ways to enhance its performance.